### v1

# A Survey of Cross-domain Graph Learning: Progress and Future Directions

Abstract
Graph learning plays a vital role in mining and analyzing complex relationships involved in graph data, which is widely used in many real-world applications like transaction networks and communication networks. Foundation models in CV and NLP have shown powerful cross-domain capabilities that are also significant in graph domains. However, existing graph learning approaches struggle with cross-domain tasks. Inspired by successes in CV and NLP, cross-domain graph learning has once again become a focal point of attention to realizing true graph foundation models. In this survey, we present a comprehensive review and analysis of existing works on cross-domain graph learning. Concretely, we first propose a new taxonomy, categorizing existing approaches based on the learned cross-domain information: structure, feature, and structure-feature mixture. Next, we systematically survey representative methods in these categories. Finally, we discuss the remaining limitations of existing studies and highlight promising avenues for future research. Relevant papers are summarized and will be consistently updated at: https://github.com/cshhzhao/Awesome-Cross-Domain-Graph-Learning.
Index Terms:
Graph Learning, Cross-domain, Structure, Feature, Structure-feature, Taxonomy.
I Introduction
Graph, or graph theory, has been playing an increasingly crucial role across numerous fields, actively contributing to various industries in the modern world, including recommendation [1, 2], finance [3, 4], weather [5], pharmaceuticals [6, 7], and communications [8, 9]. Graph data can model the structural characteristics between nodes, offering significant advantages in capturing interactions and dependencies across structures and features. Many real-world datasets—such as citation networks [10], molecular [7], and transportation networks [11]—are naturally represented as graphs. To efficiently process and analyze these complex graph data, Graph Neural Networks (GNNs) [12, 13, 14] have emerged as a leading technique. The primary goal of GNNs is to learn the relationships between nodes and their neighbors through recursive message passing and aggregation mechanisms, thereby obtaining expressive representations at the node, edge, or graph level for various downstream tasks.
Figure 1: The purpose of cross-domain graph learning aims to integrate graph knowledge across a myriad of graph domains and transfer the knowledge to other domains.
In recent years, the rapid development of foundation models, especially the large language models (LLMs), has significantly revolutionized the way people live and work, such as ChatGPT in natural language processing (NLP) and DALL-E 2 in computer vision (CV). Compared to previous deep learning models, the core advantage of foundation models lies in their ability to train on data from diverse domains, subsequently transferring the learned knowledge across different domains, encompassing three key capabilities: cross-modal, cross-domain, and cross-task [15, 16]. These three capabilities are also crucial in graph domains. For instance, cross-modal capabilities allow social networks to combine user relationships (graph structure) with textual content (text modality) to more accurately predict user interests, thus enhancing recommendation systems. Cross-domain capabilities facilitate knowledge transfer between different domains—such as leveraging insights from transportation networks to optimize power load management in power grid networks, or using rich user interaction data from social networks to improve anomaly detection in banking transaction networks with limited data. Cross-task capabilities enable models trained for one task (e.g., node classification to predict a paper’s research field) to be directly applied to another task (e.g., link prediction to identify paper citations), thus enhancing adaptability and efficiency across tasks.
The importance of cross-domain graph learning (CDGL). In graph domains, the key to achieving this vision lies in finding an effective way to integrate cross-domain knowledge that allows for straightforward generalization across domains. In detail, single-domain cross-modal problems (e.g., integrating images, text, and graphs) can benefit from advancements in CV and NLP, where many methods have successfully leveraged multi-modal information [17]. Additionally, single-domain cross-task problems can be addressed using classical graph pretraining models [18, 19, 20]. However, when dealing with cross-domain problems, the heterogeneity and diversity of graph data make it challenging for data from a single domain to capture a universal graph representation. Figure 1 shows the overview of CDGL, which can help better understand the differences across multi-domain graph data [21]:
  • (a)
    Structural differences: Graphs from different domains vary in node connectivity, with some being densely connected and others sparse. For instance, the social network among hotel staff forms a structured and top-down hierarchy with sparse connections, and the citation network among academic papers is densely interconnected.
  • (b)
    Feature differences: The dimensions and semantics of node features also differ across domains, such as social networks and molecular networks.
This diversity poses challenges for effectively integrating cross-domain graph information and producing universal cross-domain graph representations. To this end, CDGL should be considered the crucial step towards graph foundation models. Through CDGL, commonalities can be extracted from graphs across different domains, enabling the learning of graph representations with broad transferability.
[…]
Cross-domain Graph Learning
[…]
Figure 2: A taxonomy of graph models for solving cross-domain graph learning with representative examples.
Motivations. Although CDGL has increasingly received attention in graph domains—especially due to its role in constructing graph foundation models—this rapidly expanding field still lacks a systematic review, which hinders researchers from gaining a comprehensive understanding of the field and exploring future directions. Concretely, previous graph-based cross-domain methods were usually designed for specific scenarios, with limited cross-domain scope and lacking universality. Therefore, they are mainly included as subcategories in surveys within particular domains, such as graph neural network-based methods included in recommendation system surveys [1, 84].
To fill this gap, in this work, we aim to provide a systematic survey on CDGL, helping researchers better understand the current state of research and its challenges. Specifically, we go through existing representative CDGL works to organize our first-level taxonomy, as shown in Figure 2, categorizing based on the information captured by graph neural networks used for cross-domain transfer (i.e., structure-oriented, feature-oriented, and mixture-orineted) throughout the entire model pipeline. We further refine our taxonomy and introduce more granularity to the initial categories.
Contributions. The contributions of this work can be summarized from the following three aspects. (1) A structured taxonomy. A broad overview of the field is presented with a structured taxonomy that categorizes existing works into three categories (Figure 2). (2) A comprehensive review. Based on the proposed taxonomy, the current research progress of cross-domain techniques for graph learning is systematically delineated. To the best of our knowledge, this is the first comprehensive survey for CDGL. (3) Some future directions. We discuss the remaining limitations of existing works and point out possible future directions.
II Preliminary
In this section, we first give the related notations and then introduce the basic concepts of two key areas related to this survey, i.e., graph representation learning and cross-domain learning. Finally, we give a brief introduction to the newly proposed taxonomy for CDGL.
II-A Notations
Consider a graph instance represented as $G=\{V,E\}$, where $V=\{v_{1},v_{2},\ldots,v_{N}\}$ is the set of nodes with $N$ elements. The set $E\subseteq V\times V$ defines the connections between nodes. Associated with each node $v_{i}$ is a feature vector $x_{i}\in\mathbb{R}^{D}$. To describe the connectivity within the graph, we use the adjacency matrix $A\in\{0,1\}^{N\times N}$, where the matrix entry $A_{i,j}$ is defined as 1 if and only if there is an edge $(v_{i},v_{j})\in E$. A graph data space, including a node space and an edge space in a graph space, can be denoted as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ and $\mathcal{E}$ are the set of nodes and edges in the corresponding graph space.
II-B Graph Representation Learning
Definitions. Graph representation learning is a powerful technique for managing complex and heterogeneous non-Euclidean graph data. It focuses on learning meaningful representations for nodes, edges, or entire graphs by mapping them into continuous vector spaces. This approach preserves the intrinsic structural and feature information of the graph, which can facilitate various tasks such as node classification, link prediction, and graph classification. These methods can be broadly categorized into two main branches: shallow embedding methods and deep graph neural networks (GNNs).
Shallow Embedding Methods. Shallow embedding methods, such as node2vec [85] and DeepWalk [86], aim to map nodes into low-dimensional embeddings by preserving the network’s similarity structure. These methods utilize techniques like factorization-based methods or random walks to achieve this goal. While providing flexibility for various downstream tasks, shallow embedding methods face limitations: they cannot generate embeddings for unseen nodes and lack the capability to incorporate node features effectively.
Deep Graph Neural Networks (GNNs). Deep GNNs, including Graph Convolutional Networks (GCN) [12] and Graph Attention Networks (GAT) [13], maintain input node features as constants and optimize model parameters for specific tasks. These networks typically employ a message-passing mechanism, where nodes iteratively exchange and aggregate information with their neighbors. This process leads to more expressive graph representations that can effectively handle more complex graph-based tasks. In detail, the forward process of the message-passing mechanism can be defined as:
$h_{i}^{(l)}=\mathbf{U}\left(h_{i}^{(l-1)},\mathbf{M}(\{h_{i}^{(l-1)},h_{j}^{(l-1)}\mid v_{j}\in\mathcal{N}_{i}\})\right)$
where $h_{i}^{(l)}$ is the feature vector of node $v_{i}$ in the $l$-th layer, and $\mathcal{N}_{i}$ is a set of neighbor nodes of node $v_{i}$. $\mathbf{M}$ denotes the message passing function of aggregating neighbor information, $\mathbf{U}$ denotes the update function with central node feature and neighbor node features as input. By stacking multiple layers, GNNs can aggregate messages from higher-order neighbors.
The prevalent methods in graph representation learning are currently based on GNN architectures due to their enhanced ability to capture complex relational patterns in graph data. However, both shallow embedding methods and GNNs are primarily designed to learn representations within a specific graph domain, limiting their effectiveness across different domains. This issue underscores the need for research into methods that can generalize across diverse graph-based applications, aiming to enhance the learning capacity of graph representation techniques to operate effectively across varying domains.
II-C Cross-domain Learning
Before formally introducing CDGL, we first provide an overview of cross-domain learning, a pivotal research direction in deep learning that plays a crucial role in the development of foundational models such as LLMs. Specifically, we present a unified definition of cross-domain learning and summarize common cross-domain strategies, laying the groundwork for a better understanding of CDGL.
Definitions. Given the source domains $D_{S}=\{D_{S_{1}},\dots,D_{S_{m}}|D_{S_{i}}=\{X_{S_{i}},Y_{S_{i}}\},i=1,\dots,m\}$ and target domains $D_{T}=\{D_{T_{1}},\dots,D_{T_{n}}|D_{T_{j}}=\{X_{T_{j}},Y_{T_{j}}\},j=1,\dots,n\}$, where for any $i$ and $j$, $|D_{T_{j}}\cap D_{S_{i}}|\geq 0$, the objective of general cross-domain learning is to map the data from source domains to a shared embedding space and then learn a set of task functions based on this shared space for the target domains. Each dataset is referred to as a distinct domain, where a domain is characterized by a unique data distribution and task context.
First, we define the model $M$ as a source encoder that maps all $X_{S_{i}}\in D_{S}$ to a unified embedding space $E$, which can be also applied for target domains:
$M:X_{S_{i}}\rightarrow E,\ \text{for}\ i=1,\dots,m.$ (1)
Next, we define the set of task functions for the target domains $F=\{f_{T_{1}},\dots,f_{T_{n}}\}$, where each task function $f_{T_{j}}$ aims to map the embeddings $M(X_{T_{j}})$ to the corresponding task result $Y_{T_{j}}$:
$F=\{f_{T_{j}}:M(X_{T_{j}})\rightarrow Y_{T_{j}},\ \text{for}\ j=1,\dots,n\}.$ (2)
In this framework, the objective is to learn the model $M$ and the task functions $f_{T_{j}}$ such that the features across all domains are unified into a shared embedding space, allowing effective knowledge transfer across domains, which in turn improves generalization and performance on the target tasks.
The objective function for training $M$ can be defined as:
$\min_{M}\sum_{i=1}^{m}\mathbb{E}_{x\sim P_{S_{i}}(x)}\left[L_{M}(M(x),Y_{S_{i}})\right],$ (3)
where $L_{M}$ can be an unsupervised or supervised loss function (e.g., contrastive loss or classification loss), and $P_{S_{i}}(x)$ represents the distribution of the source domain data.
The objective function for learning the set of task functions $F$ can be defined as:
$\min_{f_{T_{j}}}\sum_{j=1}^{n}\mathbb{E}_{x\sim P_{T_{j}}(x)}\left[L_{T}(f_{T_{j}}(M(x)),Y_{T_{j}})\right],$ (4)
where $L_{T}$ is the appropriate loss function for the target domain tasks.
Note that these two processes can be conducted either through joint training or separate training. The conditions $|D_{T_{j}}\cap D_{S_{i}}|$ reflect, to some extent, the difficulty of transferring cross-domain knowledge during the training process.
The development of cross-domain learning in various fields. To better understand this research direction, we introduce the development of cross-domain learning in the fields of Natural Language Processing (NLP) and Computer Vision (CV), focusing on how cross-domain learning is achieved within these domains.
  • 1
    NLP. The development of cross-domain learning in NLP has progressed significantly, driven by innovations in model architectures, training strategies, and data availability. Early efforts relied on domain adaptation techniques, such as feature alignment and adversarial learning, to transfer knowledge from resource-rich domains to low-resource ones. Alongside these, domain generalization techniques, which aim to improve model robustness by learning domain-invariant representations, have also been widely explored. The rise of pre-trained language models (PLMs), like BERT and GPT, enabled generalization across domains by leveraging contextual embeddings and task-specific fine-tuning. More recently, large language models (LLMs), such as GPT-3 and GPT-4, have demonstrated strong cross-domain capabilities through massive pretraining on diverse corpora, excelling in zero-shot and few-shot learning. Furthermore, multi-modal techniques have extended cross-domain learning beyond text, allowing models to integrate information from other modalities, such as images and structured data, to address complex tasks like visual question answering and multi-modal dialogue.
  • 2
    CV. The development of cross-domain learning in computer vision (CV) follows a similar trajectory to that of NLP. Early approaches rely on domain adaptation techniques such as adversarial training, feature alignment, and style transfer to address challenges posed by domains with different visual characteristics. As the field advances, domain generalization methods, including invariant feature learning and meta-learning, emerge to improve model robustness to unseen domains. The introduction of pre-trained vision models like ResNet and Vision Transformers (ViTs) further the progress by enabling models to generalize across domains through the reuse of learned features, which can then be fine-tuned for downstream tasks. Recently, multi-modal models like CLIP and Sora demonstrate exceptional cross-domain capabilities by integrating visual and language understanding. For instance, Sora excels in tasks like image captioning, visual question answering, and zero-shot reasoning, benefiting from large-scale pretraining on multi-modal datasets. Just as in NLP, the combination of visual and textual data allows these models to achieve superior cross-domain performance.
  • 3
    Summary. The rapid development of cross-domain learning in NLP and CV can be attributed to the relatively straightforward alignment of features across domains, such as text and images. These fields benefit from shared representation formats—text and images—where semantic or visual features are easier to align across different domains. As a result, when the data scale increases, models can more easily learn shared cross-domain patterns, enabling better generalization.
In contrast, graph data presents unique challenges for cross-domain learning. Unlike text or images, graphs are not inherently structured and require manual definition of their structures, leading to complexity in graph construction and a lack of uniformity in graph representations. Additionally, graph features often vary widely across domains, and some features may not even carry semantic information, making them difficult to align. This inherent complexity and diversity in graph data contribute to the slower development of cross-domain learning in the graph domain, which remains an area of active research.
II-D Cross-domain graph learning
CDGL is a key yet underexplored area within cross-domain learning. Building on the previous definition, CDGL can be described by specifying the $X_{T_{i}}$ using graph data $\mathcal{G}_{T_{i}}$.
Definitions. Derived from cross-domain learning, CDGL aims to capture cross-domain knowledge from given source-domain graph data $\mathcal{G}=\{\mathcal{G}_{S_{1}},\dots,\mathcal{G}_{S_{i}}|i=1,\dots,m\}$ and then transfer the captured knowledge to learn a set of task functions (e.g., node classification, graph classification, etc.) on target-domain graph data $\mathcal{G}=\{\mathcal{G}_{T_{1}},\dots,\mathcal{G}_{T_{i}}|i=1,\dots,n\}$.
II-D1 The Refinement of Cross-domain Scale
The source and target domains in cross-domain graph learning (CDGL) are typically defined manually, and this definition plays a crucial role in determining the scope of cross-domain learning. The manner in which these domains are defined influences the scale, which in turn reflects the complexity and diversity of the domains involved. In this survey, we categorize CDGL into three distinct scales, as described below:
  • 1
    Limited cross-domain graph learning refers to scenarios where graph samples from source and target domains share a set of similar or consistent attributes (i.e., from the same field). For example, NaP [35] divides the Facebook100 dataset—which contains social networks from 100 U.S. universities—into different domains, treating the social network of each university as a unique domain. The data from each domain share similar attributes, except for the university types. CDGL is then conducted across different universities. Similarly, SDGA [44] treats ACMv9, Citationv1, and DBLPv7—datasets constructed from ArnetMiner [87]—as distinct domains for cross-domain learning.
  • 2
    Conditional cross-domain graph learning refers to situations where the boundaries between source domains and graph domains are clearly defined, but their attributes can be transformed into a common format (e.g., text) and covered by a specific high-level field. For example, OFA [38] can convert cross-domain graph data into a text-attribute graph (TAG) format. However, it faces difficulties when applied to non-TAG domains, such as brain networks [60].
  • 3
    Open cross-domain graph learning refers to scenarios where graph data from both source and target domains can be arbitrary and do not need to satisfy specific conditions. For example, GCC [26] focuses on leveraging only the structural information from datasets across domains, using it to pre-train models and learn common structural patterns in graph data. GCOPE [21] presents a general framework for cross-domain graph pre-training that unifies the feature dimensions of graph data with distinct boundaries, enabling pre-training across domains for better generalization. Importantly, these unified graph features can represent data from any domain.
In general, from the perspective of node features, if a CDGL method requires that the feature dimensions across datasets remain consistent, it typically falls under the limited scale; if it relies on mapping features into an existing shared representation space (e.g., a text space), it is conditional; and if it can align inherently diverse feature dimensions for cross-domain knowledge transfer, it belongs to the open scale.
II-D2 The Refinement of Cross-domain Difficulty
Beyond cross-domain scales, we further discuss the difficulty of learning cross-domain knowledge during training (i.e., cross-domain difficulty) to better understand CDGL. Within the same cross-domain scale, different methods encounter varying levels of difficulty due to differences in the overlap between source and target domains. Consequently, we delineate cross-domain difficulty into three distinct levels based on the magnitude of the intersection between the source and target domains:
  • 1
    High-Difficulty (When $|D_{T_{j}}\cap D_{S_{i}}|=0$): In this case, the source and target domains share no common data or features, indicating a complete lack of overlap. This extreme training setting is often adopted when the target domain has very limited data. It also aligns with the ultimate goal of CDGL—achieving open cross-domain graph learning under challenging conditions. A model that performs well here typically demonstrates strong generalizability.
  • 2
    Moderate-Difficulty (When $0<|D_{T_{j}}|-|D_{T_{j}}\cap D_{S_{i}}|<|D_{T_{j}}|$): In this scenario, the target domain shares either partial or complete overlap with the source domain. The difficulty is moderate because CDGL models can learn shared cross-domain knowledge in advance, reducing the overall challenge. However, unique domain characteristics may still require fine-tuning or domain-specific adjustments to address discrepancies.
  • 3
    Low-Difficulty (When $|D_{T_{j}}|-|D_{T_{j}}\cap D_{S_{i}}|=0$): This scenario indicates that the target domain is entirely equal to the source domain, with both domains sharing the same representation space. Cross-domain learning in this setting is relatively easy because the model does not need to transfer knowledge from source to target domains but only needs to capture cross-domain knowledge directly. This situation is more suitable for cases where both domains have abundant data, and the goal is to use a single model to handle tasks across different domains. Note that, in general, high overlap does not necessarily guarantee good generalization.
By addressing these levels, we can better evaluate and develop models for diverse scenarios. Note that some approaches (e.g., domain adaptation approaches), which explicitly delineate source and target domains and ensure that the datasets do not overlap, actually involve both labeled source and unlabeled target samples in training. In such cases, we should consider the source and target domains as completely overlapping, representing a low-difficulty scenario. Ultimately, achieving robust performance in high-difficulty settings is key to advancing open cross-domain graph learning.
II-D3 Proposed Taxonomy.
Considering the key elements of graph data (including nodes, edges, adjacency matrix, and node features, which collectively encompass both structural and feature information) and referring to the definition of the given CDGL, we propose a taxonomy (as illustrated in Figure 2) that organizes representative techniques focused on capturing generalizable cross-domain information into three main categories: (1) Structure-oriented approach, where the graph structural information from various domains is utilized as the basis for cross-domain learning. (2) Feature-oriented approach, where the graph feature information from various domains is specifically learned and leveraged. (3) Mixture-oriented approach, where both the graph structural and feature information from various domains are organically integrated and analyzed from a unified perspective.
In the following sections, we present a comprehensive survey along the three main categories of our taxonomy for learning cross-domain graph knowledge in graph-related tasks, respectively.
III Structure-oriented Approach
Structural information, as the distinctive and core characteristic of graph data, is essential for addressing real-world graph-related problems. Existing graph models have demonstrated the ability to effectively capture structural information in single-domain scenarios. However, in cross-domain scenarios, the graphs from different domains exhibit varying node connectivity patterns, with some being densely connected and others more sparsely connected. These structural differences pose significant challenges to current graph models. To address this, structure-oriented approaches aim to improve the performance of existing models by extracting commonalities in structural information across domains. Based on how structural information is utilized, these approaches can be broadly categorized into three main branches: Structure Generation and Structure Contrast.
Figure 3: The core ideas behind structure-oriented approaches for capturing cross-domain knowledge consist of two main strategies: a) structure generation approaches, which employ a shared generator module to generate new and distinct structures for graph data from both source and target domains; b) structure contrast approaches, which construct positive and negative sample pairs for various graph data or leverage graph structure information to train GNNs and then make model parameter contrast across domains.
III-A Structure Generation
To effectively capture structural information, a useful approach is to establish a unified node connection strategy for graphs, ensuring that the graph data in both source domains and target domains follow the same connection rules. This allows the graph data from different domains to share common structural characteristics. The CDGL method based on structure generation is proposed with this idea in mind. It employs a shared module to generate new and distinct structures for graph data from both source and target domains. These newly generated structures enable the graph model to adopt a unified perspective for analyzing structural information across various domains, thereby enhancing the model’s performance on different graph-related tasks, as shown in Figure 3. For instance, GraphControl [22] introduces ControlNet, designed specifically for graph data, to extract downstream-specific graph structures from target domain graphs. Finally, by incorporating relevant feature information into the extracted subgraph data, cross-domain learning is achieved through fine-tuning or prompting. UniAug [25] is the first to introduce diffusion models for CDGL. Training a graph diffusion model on multi-source graph structure data enables the generation of transferable graph structures tailored to various downstream domains. GA2[23] is the first to reformulate graph data from different domains into subgraphs. In the pretraining phase, it leverages an autoencoder architecture to embed and reconstruct these subgraphs, enabling the graph model to utilize diverse structural information. OpenGraph [24] carefully a structure-based node tokenization scheme for graph domains, then introduces a scalable graph transformer to learn the node token embeddings and subsequently employs the LLM techniques to generate different graph structures, finally tackle various downstream graphs with variations in node sets and feature semantics.
III-B Structure Contrast
In contrast to structure generation, the goal of structure contrast is to effectively compare the structures of different graph datasets and identify common structural patterns across domains, as shown in Figure 3. This approach centers on designing negative and positive sample pairs to enable meaningful comparisons between graph structures, or on learning a set of weights to compare different structure-related model parameters. For instance, GCC [26] employs structural contrastive learning to pre-train graph neural networks in a cross-domain manner, enabling models to learn transferable structural representations across diverse domains. PCRec [27] pre-trains a graph neural network using self-supervised contrastive learning on source domain graphs to capture transferable structural information, then transfers the pre-trained encoder to initialize node embeddings on the target domain, enhancing recommendation performance by balancing source and target information while reducing bias from source domains. GRADE [28] introduces a graph adaptive network for CDGL by measuring and minimizing structural distribution shifts between source and target graphs using a Graph Subtree Discrepancy based on the Weisfeiler-Lehman graph isomorphism test, effectively comparing structural information for improved transferability. FedStar [30] leverages federated learning to train structure-only graph models on different clients, compares their parameters, and integrates them via weighted averaging to create a global model that merges cross-domain structural information. APT [29] leverages structural properties to contrast graph data during cross-domain pretraining, developing graph selection strategies that help the pretrained model more effectively capture cross-domain structural patterns. BooG [31] boosts the cross-domain learning of graph foundation models by unifying structural characteristics through virtual super nodes and edges, and employs a novel contrastive learning pre-training objective to learn expressive, generalizable graph representations. ProCom [32] enhances cross-domain targeted community detection by capturing structural information of latent communities through a dual-level context-aware pre-training method and a targeted community-guided prompting mechanism. RiemannGFM [33] captures cross-domain knowledge by learning a universal structural vocabulary of graph substructures—such as trees and cycles—embedding them into Riemannian manifolds and leveraging geometric contrastive learning for robust and transferable representations.
III-C Discussions
Structure-oriented approaches have demonstrated superior performance on realizing CDGL, being able to handle various complex graph structural information. Moreover, they, specifically the structure generation approaches, also exhibit strong flexibility, as the unified node connection rules can generate meaningful structures for the graphs from unseen domains. Another advantage of these methods is their ability to explore the potential for uncovering commonalities across graph data from different domains, with structural information serving as the foundation of such data. For graph data, while structural information forms the foundation, it is the feature information that plays a crucial role in conveying substantive meaning. However, structure-oriented approaches mainly focus on capturing structural information but sometimes ignore the feature information. Taking structure generation approaches as an example, they need to consider the coherence between the generated structure and feature information to ensure effective communication across graph data from different domains.
III-D Significance and Future Implications
Existing structure-oriented CDGL methods hold significant potential for bridging diverse domains by capturing universal structural insights. They effectively uncover structural commonalities across domains and achieve generalization at an open scale. Although relying solely on structural information may limit performance, these approaches provide critical insights into transferring structural patterns, paving the way for open CDGL and graph foundation models that seamlessly integrate structural and feature information.
Figure 4: The core ideas of feature-oriented approaches: a) feature align approaches, which directly leverage a trainable feature extractor to mine the commonalities among source and target domains; b) feature misalign approaches, which need to align the graph data with different feature dimensions across domains for training a general graph models. Note that both approaches can be further enhanced by prompting techniques.
IV Feature-oriented Approach
As discussed in Section III-C, feature information plays a vital role in conveying substantive meaning in graph data, helping to differentiate between domains. The essence of feature-oriented approaches is to integrate feature information from various domains to discover commonalities across cross-domain features. However, training a general graph model on data from multiple domains presents unique challenges, primarily due to the feature differences across domains, particularly when feature dimensions do not align. For example, in most cases, if the feature dimensions of source domain $A$ and target domain $B$ differ, their semantic meanings will inevitably differ as well, since it is highly unlikely that domains with different dimensions share the same semantics. On the other hand, when A and B have the same feature dimensions, the semantics may either be identical or distinct. If the semantics are aligned, no additional adjustments are needed, and the domains can essentially be treated as one. However, when the semantics differ, further steps are required to bridge the semantic gap. Therefore, in this section, we broadly categorize models into Dimension Align and Dimension Misalign approaches, based on whether they require feature dimension alignment.
IV-A Dimension Align
The earliest feature-oriented approaches are Dimension Align, meaning they perform CDGL when the feature dimensions of the graph data in the source and target domains are consistent. Typically, when the feature dimensions align, the source and target domains share a high-level domain. Dimension Align approaches are generally considered part of limited CDGL. For instance, KTN [34] introduces different feature extractors for different node types in heterogeneous graph neural networks, and then uses domain adaptation strategies to integrate cross-domain feature information. NaP [35] improves out-of-distribution generalization in graph contrastive learning by dynamically treating semantically similar cross-domain negative pairs as positive pairs based on embedding similarity, reducing the distribution gap between domains, achieving CDGL to some extent. GPF-plus [36] and RELIEF [37] are two newly proposed feature-based graph prompt techniques to align the pre-training knowledge to different downstream datasets, achieving limited CDGL across molecular domains, where the dimension of different molecular datasets can be represented in the same feature space [88].
IV-B Dimension Misalign
Since Dimension Align approaches are significantly limited in cross-domain scale compared to open CDGL, which has already been widely achieved in NLP and CV thanks to the rapid development of LLMs, Dimension Misalign approaches have been proposed in recent years. These methods can map graph features from different domains into a shared embedding space with the same dimension. By aligning graph feature dimensions, they relax the limitation of cross-domain scale. Depending on the feature mapping strategy used, these approaches are classified as either conditional or open CDGL. For example, OFA [38] and ZeroG [41] both use LLMs to align text-attributed features from different domains. OFA freezes LLMs and directly trains a Multi-layer Perceptron (MLP) on aligned node features to achieve zero-shot node classification, while ZeroG [41] further utilizes a LoRA fine-tuning strategy to align semantic information across different domains and then leverages a text similarity task to conduct node classification results, achieving CDGL. Similar to the former approaches, GraphAlign [39] also uses LLMs to algin features, while it additionally designs alignment strategies of feature encoding, normalization, alongside a mixture-of-feature-expert module. Besides, CDFS-GAD [40] employs a domain-adaptive graph contrastive learning module that enhances cross-domain feature alignment by optimizing contrastive loss to bring similar nodes closer and push dissimilar ones apart. Additionally, a domain-specific prompt-tuning module customizes the encoder to effectively capture and retain both common and unique features across domains. Furthermore, DAGPrompT [42] proposes GLoRA to adapt feature projection through low-rank matrix updates, aligning pre-trained GNNs with downstream task distributions.
IV-C Discussions
Compared with structure-oriented approaches, focusing on learning cross-domain feature commonalities to build graph models also demonstrates its own distinct advantages in handling CDGL especially when the feature information between the source domains and target domains differs significantly. However, because feature-oriented approaches often skip the dedicated extraction of structural information, they may encounter conflicts between structural and feature patterns when training a cross-domain graph model, limiting the performance upper bound.
IV-D Significance and Future Implications
If structure-oriented CDGL research serves as the basis for achieving open CDGL or graph foundation models, then feature-oriented CDGL represents a critical breakthrough toward realizing true open cross-domain scalability. As mentioned earlier, the feature dimensions of graph data from different domains are often inconsistent, and even when they align, their semantic meanings may differ. This makes it exceedingly challenging to learn commonalities from cross-domain features. Existing feature-oriented CDGL works extensively explore how to comprehend feature information across domains and effectively capture their shared semantics. These insights are pivotal for advancing toward open CDGL, making feature-oriented approaches a key enabler of this vision.
Figure 5: The core ideas of mixture-oriented approaches. a) Feature-Structure Mixture: Feature information is prioritized, and structure information is integrated afterward. b) Structure-Feature Mixture: Structure information is prioritized, and feature information is integrated afterward. c) GNN-based Unified Mixture: Structural and feature information are jointly extracted and processed using a unified extractor during GNN training. d) Flatten-based Unified Mixture: Graph data is transformed into textual sequences through a flattening process, enabling large language models (LLMs) to handle both structural and feature information for downstream tasks. This categorization details two distinct integration philosophies: Sequential-Mixture (a-b) and Unified-Mixture (c-d).
V Mixture-oriented Approach
Mixture-oriented approaches are capable of performing cross-domain learning on both structural and feature information simultaneously, exploring the commonalities of structure and feature either individually or in pairs. This has been the most mainstream research direction in CDGL. Based on the different ways of combining structure and feature information, we categorize Mixture-oriented approaches into two types: Sequential Mixture and Unified Mixture.
V-A Sequential Mixture
Similar to sequential ensemble learning techniques [89, 90], the Sequential Mixture approach typically processes structural information and feature information in a sequential manner. This can involve first processing the structural information, followed by the introduction of feature information. For instance, SDGA [44] first leverages a graph convolutional network (GCN) to extract high-order structural information, followed by an adversarial transformation module with shift parameters to align the source graph feature distribution with that of the target domain, thereby achieving sequential mixture. TAPE [45] leverages LLMs to generate textual explanations as enriched features, which are then used by GNNs to integrate structural information, ultimately enhancing the performance of node classification on text-attributed graphs. GSPT [47] proposes to construct context using random walks first, which is used to represent structural information of the graph, and then feed the context into a transformer to perform masked feature reconstruction, which means the feature information is reconstructed afterward. SOGA [46] first leverages the structural properties of the target graph through Structure Consistency optimization before enhancing feature information through Information Maximization, enabling effective adaptation to the target domain without access to source data. Alternatively, the method can prioritize the feature information before incorporating the structural information. An example of this is UniGraph [49] starts with textual features (raw text features associated with nodes) being fed to the language model encoder (DeBERTa, E5). After processing the textual features, the resulting embeddings are then passed to the GNN encoder for further propagation across the graph. PTGB [43] prioritizes feature transformation and alignment through a data-driven atlas mapping approach before capturing structural properties during GNN pre-training. ARC [48] aligns features across different graph datasets into a common anomaly-sensitive space before utilizing residual-based graph encoding to capture high-order structural information for anomaly detection. AnyGraph [50] employs a feature-centric pre-processing approach using SVD and feature mapping before handling graph structures through an expert routing mechanism and high-order connectivity encoding, allowing for effective cross-domain adaptation. GraphLoRA [51] leverages a structure-aware contrastive low-rank adaptation module that integrates structural information from the target graph, mitigating structural discrepancies, while its Structure-aware Maximum Mean Discrepancy (SMMD) aligns node feature distributions across graphs, ensuring both structural and feature-level adaptation. SAMGPT [52] first aligns and adapts features across domains, then learns universal structure tokens during multi-domain pre-training, and finally utilizes holistic and specific prompts to transfer both unified and domain-specific structural knowledge to downstream tasks. MDGFM [53] first projects domain-specific features into a unified semantic space using unique domain tokens, and then aligns graph topologies through a graph topology-aware alignment mechanism that, via contrastive learning, extracts universal structure tokens for effective prompt-based adaptation.
V-B Unified Mixture
In contrast to the Sequential Mixture approach, the Unified Mixture approach typically considers structural and feature information from a unified perspective during the training process, with the goal of achieving an organic mixture. As shown in Figure 5, unified mixture approaches typically involve two steps: (1) employing a unify function $Unify(\cdot)$ to extract graph structures and features simultaneously to generate node/edge/graph-level representations $H_{rep}$, and (2) a task head function $TaskH(\cdot)$ is then leveraged to mapping the representations and output the predicted results, as illustrated below:
Mixture Unifying: $H_{rep}=Unify(\mathcal{V},\mathcal{E},\mathcal{A},\mathcal{X}),$ (5)
Prediction: $\hat{Y}=TaskH(H_{rep}),$ (6)
where $\mathcal{V},\mathcal{E},\mathcal{X}$ denotes the set of nodes, edges, and node features, respectively. $\mathcal{A}$ is the adjacency matrix of the given graph.
Traditionally, the strategies for $Unify(\cdot)$ and $TaskH(\cdot)$ are part of a cohesive framework. The $Unify(\cdot)$ strategy is typically built upon the message-passing mechanism, where the differences lie in the ways of passing or the applied information encoding techniques. Subsequently, $TaskH(\cdot)$ is realized through a GNN-based task head, such as fully connected layers, which then outputs the task prediction results based on the $H_{rep}$ from $Unify(\cdot)$. However, with the advent of the powerful representation and reasoning capabilities of LLMs in NLP, flatten-based unifying strategies have been introduced into graph learning. These strategies can fundamentally transform graph data into semantically rich node or token sequences. LLMs then leverage these sequences to perform reasoning and generate corresponding task results [91]. To help comprehensively understand these unified mixture approaches, the following sections will illustrate these two sets of unify-task strategies in detail.
GNN-based Unified
In general, GNN-based unified mixture approaches usually adopt similar selection strategies of task head functions while differing according to the passing ways or the applied information encoding techniques. For example, DA-GCN [54], UDA-GCN [54], COMMANDER [56], CCDR [61], GAST [62], DH-GAT [63], ALEX [64], ACT [65], ALCDR [69], and DGASN [67] all adopt domain adaptation training frameworks, differing mainly in their approaches to unifying data from source and target domains. Specifically, DA-GCN and UDA-GCN utilize graph convolutional networks (GCNs) with domain adaptation, which explores the initial possibilities for Unified Mixture CDGL by employing GCNs to aggregate structural and feature information together in a preliminary unified manner. To better handle the unified structure and feature information, more advanced methods have been introduced, such as graph attention mechanisms, hyperbolic representations, and contrastive learning. COMMANDER employs a graph-attentive encoder for domain alignment, providing more flexibility and dynamic focus during feature aggregation. CCDR proposes a unified contrastive learning framework for cross-domain recommendation, utilizing both intra-domain and inter-domain contrastive learning to enhance representation learning and effectively transfer user preferences between domains. GAST uses a hybrid graph attention mechanism along with a POS-Transformer to integrate syntactic and sequential semantics, enhancing both structural and semantic representation. DH-GAT applies hyperbolic graph attention networks to model hierarchical structures effectively, leveraging the power of hyperbolic space to represent latent relationships. ALEX leverages singular value decomposition (SVD) and domain discriminators to unify cross-domain representations, facilitating effective feature alignment. ACT employs contrastive learning with domain alignment for anomaly detection, focusing on learning similarities and differences between domains to generate cohesive representations. ALCDR uses an optimal transport-based anchor link learning mechanism to adaptively transfer knowledge across domains, providing a robust means of cross-domain unification. DGASN employs a graph attention network (GAT) to simultaneously learn node and edge embeddings, and using adversarial domain adaptation to align the source and target distributions, thus achieving domain-invariant graph representations.
Additionally, CDGEncoder [58], PGPRec [70], STGP [68], GCOPE [21], UniGLM [73], CrossLink [71], and MDGPT [72] all involve a pretraining phase, but they differ in their information encoding or propagation methods during pretraining. Specifically, the CDGEncoder uses a multi-view GNN encoder combined with attention mechanisms to aggregate contextual and topological views. PGPRec introduces personalized graph prompts and uses contrastive learning during pretraining to enhance user/item representations. STGP utilizes a two-stage prompting strategy to adapt both domain and task-specific properties across spatio-temporal tasks. GCOPE employs graph coordinators to align diverse graph datasets during pretraining, mitigating negative transfer. UniGLM considers introducing LLMs as the encoder for text attributes in the traditional pre-training paradigm with a tailored contrastive learning approach based on the adaptive positive sample selection to align features and structures. CrossLink leverages a decode-only transformer to model the temporal evolution of dynamic graphs for link prediction. Finally, MDGPT uses domain tokens and dual prompts to unify and adapt multi-domain graph features, enhancing cross-domain generalization.
In addition to the above two mainstream paradigms, domain generalization, multi-task learning, and federated learning have also been explored. For example, CrossHG-Meta [59] proposes to process structural and feature information simultaneously by aggregating heterogeneous information from multiple semantic contexts and then designs a unified domain generalization framework for CDGL, especially for heterogeneous graphs. METABrainC [60] proposes a pre-training cross-domain framework via meta-learning techniques using multiple self-supervised tasks on source datasets and adapts it to target tasks via fine-tuning, enhancing cross-domain brain connectome analysis. GCFL [57] employs a clustering algorithm to group graph models trained on clients based on parameter similarities, thereby achieving a unified representation of structural and feature information within a federated learning framework, ultimately enhancing cross-domain generalization.
Flatten-based Unified
Referring to the input of LLMs, flatten-based Unified Mixture approaches focus on converting graphs into textual descriptions by transforming them into sequences of nodes or tokens. These approaches then develop tailored learning strategies to allow LLMs to process these sequences and generate task-specific outputs effectively. For instance, GIMLET [74] introduces a unified graph-text approach called Generalized Position Embedding, which transforms graphs into token sequences and seamlessly combines graph structure with LLM-driven text processing. GraphTranslator [75] employs a Translator module to convert node embeddings into token sequences, enabling effective integration of pre-trained graph models and LLMs, allowing LLMs to handle both pre-defined and open-ended graph tasks. GraphGPT [76] leverages a dual-stage instruction tuning mechanism to align graph structural and semantic information with LLMs, enhancing cross-domain generalization through both graph structural alignment and node feature transfer. GITA [77] combines graph visualization with textual conversion, using a Vision-Language Model (VLM) to jointly reason over visual and textual graph data, providing a comprehensive integration of graph semantics and structure. InstructGraph [78] uses a structured format verbalizer to convert graph data into text sequences, combined with instruction tuning and preference optimization to improve graph reasoning and reduce hallucinations, making LLMs suitable for graph-centric tasks. TEA-GLM [79] employs a linear projection module to align pre-trained Graph Neural Network (GNN) representations with LLM token embeddings, enabling zero-shot predictions across multiple graph tasks without modifying the LLM itself. GOFA [80] introduces a hybrid architecture that interleaves GNN layers with a pre-trained LLM, combining structural and semantic modeling abilities while transforming graph data into flattened text descriptions for LLM-based text generation. HiGPT [81] utilizes a heterogeneous graph tokenizer to transform complex graph structures into textual representations, employing a novel instruction-tuning framework to enhance LLM’s understanding of intra- and inter-node relationships. GraphCLIP [82] uses a graph-summary contrastive learning approach by generating textual descriptions of subgraphs with an LLM, which improves zero-shot and few-shot transfer capabilities through graph prompt tuning and invariant learning. GraphWiz [83] leverages the GraphInstruct dataset to train LLMs in solving graph computational problems, using an instruction-following framework that transforms graphs into textual sequences, allowing explicit reasoning paths for improved task performance.
V-C Discussions
Mixture-oriented approaches mainly focus on simultaneously integrate structure and feature information across domains, enabling deep learning models to more comprehensively capture commonalities across domains. Nevertheless, the absence of a powerful feature alignment technique that effectively maps datasets from different domains into a shared feature space remains a significant limitation for feature-oriented approaches. As shown in Table I, earlier mixture-oriented approaches were only able to scale to limited cross-domain settings. Recently, some feature alignment techniques, such as large language models (LLMs), have been introduced, but they have only managed to achieve performance at a conditional scale. Indeed, a few newly proposed methods have demonstrated the potential to achieve open cross-domain scale, but their primary focus has been on exploring the feasibility of true open CDGL. For example, some approaches utilize domain-specific linear projection layers to align feature dimensions across domains, yet the resulting performance improvements have been limited.
V-D Significance and Future Implications
Mixture-oriented approaches are fundamentally designed to effectively integrate structure-oriented and feature-oriented methods, aiming to achieve a synergistic effect where the whole is greater than the sum of its parts (i.e., $1+1>2$). For instance, since the cross-domain representation spaces derived from structural and feature information are inherently distinct, finding an organic way for these spaces to adapt to each other is both challenging and crucial. Current explorations in mixture-oriented approaches not only offer valuable insights but also pave the way for achieving truly open CDGL and even the development of graph foundation models.
Model Backbone Paradigm Scale Difficulty LLM Fine-tuning Prompting Domain Setting Task Code
Structure-oriented GCC[26] GIN Pre-training Open High - Social Networks, Academic Networks, etc Full-shot Node, Graph Link
PCRec[27] GIN Pre-training Limited High - Recommendation Full-shot Link -
APT[29] GIN Pre-training Open High - Citation Networks, Protein, etc Full-shot Node, Graph Link
GRADE[28] GCN Domain Adaptation Open Low - Citation Networks, Recommendation, etc Unsupervised Node, Link Link
FedStar[30] GCN, GIN Federated Learning Open Low - Bioinformatics, Social Networks, etc Full-shot Graph Link
ProCom[32] GCN Pre-training Conditional High - Social Networks, E-commerce, etc Few-shot Community Detection Link
GraphControl[22] GIN Pre-training Open High - Social Networks, Academic Networks, etc Full/Few-shot Node, Graph Link
GA2E[23] GAE Pre-training Open Moderate - Citation Networks, Community Networks, etc Full-shot Node, Link, Graph -
BooG[31] GAT, GIN Pre-training Conditional High Sentence Transformer Citation Networks, Molecular, etc Zero/Few-shot Node, Link, Graph Link
UniAug[25] GCN, GIN Pre-training Open Moderate - Biological Networks, Chemical Networks, etc Full-shot Node, Link, Graph Link
OpenGraph[24] GT Pre-training Open High - Recommendation, Social Networks, etc Zero/Few-shot Node,Link Link
RiemannGFM[33] - Pre-training Open High - Citation Networks, Social Networks, etc Full/Few-shot Node,Link Link
Feature-oriented KTN[34] HGNN Domain Adaptation Limited Low - Academic Networks Zero-shot Node -
NaP[35] GCN Domain Generalization Limited High - Social Networks Full-shot Node -
CDFS-GAD[40] GraphSAGE Domain Adaptation Open Low - Social Networks Few-shot Anomaly Detection -
OFA[38] GCN, GAT, etc Pre-training Conditional High Sentence Transformer, etc Citation Networks, Molecular, etc Zero/Few-shot Node, Link, Graph link
ZeroG[41] GCN, GAT, etc Pre-training Conditional High Sentence Transformer, etc Citation Networks Zero-shot Node link
GraphAlign[39] GCN Pre-training Open High E5-small Citation Networks, Knowledge Graphs, etc Few-shot Node, Link link
RELIEF[37] GCN Pre-training Limited High - Citation Networks, Molecular, etc Few-shot Node, Graph link
GPF-plus[36] GIN Pre-training Limited Moderate - Citation Networks, Knowledge Graphs, etc Few-shot Node, Link link
DAGPrompT[42] GCN LoRA Limited High - Citation Networks, Molecular, etc Full/Few-shot Node,Graph Link
Mixture-oriented UDA-GCN[54] GCN Domain Adaptation Limited Low - Citation Networks Full-shot Node -
COMMANDER[56] GAT Domain Adaptation Limited Low - - Yelp, Amazon Unsupervised Anomaly Detection -
DA-GCN[92] DA-GCN Domain Adaptation Limited Low - E-commerce, Video Streaming Full-shot Node -
GCFL[57] GIN Federated Learning Open Low - Molecular, Social Networks Full-shot Graph -
METABrainC[60] GCN, GAT, etc Multi-task learning Limited High - PPMI, HIV, BP Few-shot Graph -
GAST[62] HGAT Domain Adaptation Limited Low GloVe, BERT Product Full-shot Sentiment classification -
CDGEncoder[58] GIN Domain Adaptation Open Low - Molecule, Bioinformatics, etc Few-shot Graph -
CrossHG-Meta[59] - Domain Generalization Limited High - Academic Networks Few-shot Node Link
CCDR[61] GAT Contrastive Learning Conditional High - Product Zero-shot Link Link
PTGB[43] GCN, GAT, etc Pre-training Limited High - Brain Network Full-shot Node, Graph Link
SDGA[44] GCN Domain Adaptation Limited Low - Citation Networks Semi-supervised Node -
ALCDR[69] LightGCN Domain Adaptation Conditional Low - Recommendation Full-shot Link -
DGASN[67] GAT Domain Adaptation Limited Low - Citation Networks Full-shot Link Link
CDTC[66] GIN Domain Adaptation Limited Low - Molecular Few-shot Graph -
ACT[65] GraphSAGE Domain Adaptation Limited Low - Social Networks Full/Few-shot Graph Link
DH-GAT[63] GAT Domain Adaptation Limited Low BERT Newswire, Social Media Full-shot Entity Recognition -
ALEX[64] GCN Domain Adaption Conditional Low - Citation Networks, Academic Networks Full-shot Node -
PGPRec[70] GCN, GAT, etc Pre-training Conditional High - Recommendation Full-shot Link -
TAPE[45] GCN, etc Pre-training Conditional High - Citation Full-shot Node -
CrossLink[71] DyGFomer Domain Generalization Open High - Social Networks, Question Answered, etc Full-shot Link -
GCOPE[21] GCN, FAGCN Pre-training Open High - Citation Networks, Social Networks Few-shot Node, Graph Link
Uni-GLM[73] GCN, etc Pre-training Conditional Moderate BERT Citation Networks, E-commerce Networks Semi-supervised, Few-shot Node, Link Link
SOGA[46] GCN Domain Adaption Limited Low - Citation Networks Unsupervised Node -
STGP[68] STGNN Pre-training Open High - METR-LA, etc Few-shot Node link
GSPT[47] Transformer Pre-training Conditional Moderate SentenceBERT Citation Networks Few-shot Node, Link Link
ARC[48] - Pre-training Conditional High - Citation Networks, Social Networks, etc Few-shot graph anomaly detection link
UniGraph[49] GAT Pre-training Conditional High Llama, vicuna-7B, etc Citation, Molecular, etc Zero/Few-shot Node, Edge, Graph Link
TEA-GLM[79] GraphSAGE Pre-training Conditional High Vicuna-7B Citation network, e-commerce Zero-shot Node, Link link
GOFA[80] Transformer Pre-training Conditional High Llama2, Mistral, etc Citation and product Networks Zero-shot Node, Link Link
HiGPT[81] HGNN Pre-training Conditional High Sentence-BERT Movie, Academic Networks Few-shot/Zero-shot Node Link
MDGPT[72] GIN Pre-training Open High - Citation, E-commerce, etc Few-shot Node, Graph -
AnyGraph[50] GCN Pre-training Open High - E-commerce, Academic Networks, etc Zero-shot Node, Link, Graph Link
GIMLET[74] Transformer Pre-training Limited High T5 Molecular Few-shot/Zero-shot Graph Link
GraphTranslator[75] GraphSAGE Pre-traing Conditional High ChatGLM2-6B E-commerce, Academic Few-shot/Zero-shot Node Link
GraphGPT[76] Transformer, etc Instruction tuning Conditional High Vicuna-7B-v1.1/1.5 Citation Networks Zero-shot Node, Link Link
GITA[77] GCN, GAT Instruction tuning Open High GPT-4V, LLaVA-7B/13B Collaboration Networks, Citation, etc Zero-shot/Full-shot Node, Link Link
InstructGraph[78] - Instruction tuning Conditional High LLaMA2 GSM8K, NLGraph, etc Full/Few-shot Node, Link, Reasoning, etc Link
GraphCLIP[82] GCN, GAT Pre-traning Conditional High Qwen2-72B Academic,E-commerce,etc Few/Zero-shot Node, Link, Graph Link
GraphWiz[83] - Instruction tuning Limited High LLaMA2-7B/13B, Mistral-7B GraphInstruct, NLGraph, etc Zero-shot/Few-shot Reasoning Link
GraphLoRA[51] GAT LoRA Open High - e-commerce, Social Networks, etc Full/Few-shot Node Link
MDGFM[53] GCN Pre-training Open High - Citation Networks, Social Networks, etc Few-shot Node -
SAMGPT[52] GCN Pre-training Open High - Citation Networks, Social Networks, etc Few-shot Node,Graph Link
TABLE I: A summary of models that leverage cross-domain knowledge to enhance graph-related tasks in the literature, ordered by their release time. Paradigm denotes the core technique used to achieve Cross-Domain Graph Learning. Domain indicates the domains across which knowledge is transferred. Scale and Difficulty refer to the cross-domain scales and the level of difficulty. Fine-tuning denotes whether it is necessary to fine-tune the parameters of Graph Neural Networks (GNNs). Prompting indicates the use of graph- or text-formatted prompts. Acronyms in Task: Node refers to node-level tasks; Link refers to link-level tasks; Graph refers to graph-level tasks; Reasoning refers to Graph Reasoning.
VI Future Directions
Table I summarizes the graph models that transfer particular graph information to achieve CDGL according to the proposed taxonomy. Based on the review and analysis illustrated above, we believe that there is still much space for further enhancement in this field, especially for realizing the true graph foundation models. In this section, we discuss the remaining limitations of capturing graph information across domains used for achieving effective CDGL and list some directions for further exploration in subsequent research.
VI-A Feature Alignment.
The correlation between feature dimensions and semantics across graph data from different domains is a crucial aspect to consider in CDGL. Misaligned feature dimensions will hinder the training of a general graph model capable of handling graph data from various domains, while misaligned semantics will prevent the trained graph model from capturing shared characteristics across domains, thereby limiting the performance of CDGL.
In existing CDGL methods, those targeting limited-scale scenarios can often skip this issue, as the graph data from source and target domains typically share identical feature dimensions by nature, such as in molecular networks [17, 36]. For conditional scenarios, which mainly focus on graph data in recommendation systems [70] and text-attributed graphs (TAGs) [38, 41], feature alignment challenges are circumvented in different ways: Graph data in recommendation systems may come from different domains, but they are all built on the basis of users and items, thus avoiding feature alignment problems. For TAGs, features can be easily aligned using embeddings generated by large language models (LLMs).
However, these methods are insufficient for addressing open CDGL, where feature dimensions and semantics are highly complex and heterogeneous across domains. Although some methods have begun exploring feature alignment solutions for more open scenarios, the research is still in its early stages. For instance, methods like ProG [93], MDGPT [72], and CDFS-GAD [40] have attempted to align feature dimensions across different datasets using techniques such as SVD or MLP. However, even after aligning feature dimensions, the aligned features may still lack semantic equivalence. Therefore, proposing a universal feature alignment approach to handle datasets from different domains is a key research direction towards achieving effective open CDGL.
VI-B Dealing with the Scale of Cross-domain Datasets.
In NLP and CV, the cross-domain generalization ability of foundation models follows the Scaling Law, which suggests that generalization improves with an increase in model size and the number of training samples. Typically, training an effective foundation model requires a large volume of samples, often in the order of billions, whereas traditional CDGL models generally rely on much smaller training datasets. This discrepancy in data volume limits the cross-domain performance and generalization capabilities of these models. While structure-oriented methods can process large-scale graph-structured data and achieve open CDGL at the structural level, the lack of feature information restricts their performance. However, when both structural and feature information are considered, traditional CDGL methods usually achieve only limited or conditional cross-domain learning, inherently constraining the diversity and scale of the datasets used for training. Although certain models, such as GCOPE, can alleviate some limitations on dataset diversity and scale to achieve open CDGL, the scale and diversity of the datasets remain limited. Therefore, to significantly improve the performance of graph models in open cross-domain scenarios and contribute to true graph foundation models, large-scale and diverse datasets are essential.
VI-C Overcoming data leakage.
Table I demonstrates that LLMs are receiving increasing attention in CDGL, exploring their potential capabilities in this area. However, the introduction of LLMs inevitably brings about the issue of data leakage, which has become a focal point of discussion [94]. Specifically, since LLMs are pre-trained on extensive text corpora, it’s likely that they may have seen and memorized at least part of the test data from common benchmark datasets, especially in citation networks. This undermines the reliability of current studies that rely on earlier benchmark datasets. Moreover, [95] demonstrates that specific prompts could potentially enhance the ”activation” of LLMs’ corresponding memory, thereby influencing evaluation outcomes. Both [96] and [45] have attempted to avoid the data leakage issue by collecting a new citation dataset, ensuring that the test papers are sampled from time periods after ChatGPT’s data cut-off. However, these efforts remain limited to the citation domain, and the impact of graph structures in their datasets is not significant. Therefore, it’s crucial to reconsider the methods employed to accurately evaluate the performance of LLMs on graph-related tasks. A fair, systematic, and comprehensive benchmark is also needed.
VI-D Improving interpretability.
Interpretability, also known as explainability, refers to the ability to explain or present a model’s behavior in terms that humans can understand [91]. In CDGL, most methods utilize Graph Neural Networks (GNNs). However, GNNs are typically treated as black boxes [97] and lack interpretability. While these methods consider the relationships among various pieces of information in the graph during their design and have experimentally demonstrated effectiveness in CDGL, their improvements are often based on intuition and a high-level understanding of graph data. Consequently, their interpretability remains poor. To enhance interpretability, CDGL approaches that integrate Large Language Models (LLMs) have been proposed. This is primarily due to LLMs’ reasoning and explanatory abilities, which enable them to produce user-friendly explanations for graph reasoning tasks [91, 83]. However, most of these approaches are limited by their reliance on Text-Attributed Graphs (TAGs); they can only provide explanations for conditional CDGL scenarios. Extending the powerful reasoning and explanatory abilities of LLMs from TAGs to non-TAGs is also an important direction. This expansion directly affects the use of CDGL models in applications with strict security requirements, such as network fault diagnosis [9, 8].
VI-E Unified Cross-domain Evaluation Metrics.
As discussed in Sections II-C and II-D, the ultimate goal of CDGL is to develop a foundational model capable of generalizing across any domain. However, most existing approaches evaluate model performance using classical graph-related task metrics, assessing cross-domain generalization ability based on performance across different datasets. This approach may not fully capture the model’s ability to generalize from a global perspective. Consequently, designing a unified cross-domain evaluation metric to holistically assess generalization ability represents a promising direction for future research.
VI-F Compatibility Analysis and Improvement across Domains.
Existing CDGL approaches often treat data from different domains uniformly, aiming to uncover correlations in structure, features, or structural-feature patterns to enhance cross-domain graph representation learning for various tasks. However, in practice, the compatibility of information between domains varies significantly. Similar to how humans find it easier to transfer knowledge between closely related fields (e.g., applying mathematical principles to physics) than between unrelated fields (e.g., applying musical theory to physics), analyzing domain compatibility and making targeted adjustments during cross-domain learning is a critical area for further investigation.
VII Conclusion
The mining of commonalities in data from different domains has led CDGL to emerge as a prominent area of research in recent years, especially as it is a key point for achieving true graph foundation models. In this survey, we aim to provide an in-depth overview of existing strategies for transferring knowledge across domains. First, we introduce a novel taxonomy that categorizes techniques involving various cross-domain scales into three categories based on the different graph information transferred by specific models: structure, features, and a mixture of both. Next, we systematically review representative studies according to this taxonomy. Finally, we discuss some limitations and highlight several future research directions. Through this comprehensive review, we aspire to shed light on the advancements and challenges in the field of CDGL, thereby encouraging further progress in this domain.
Acknowledgments
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.